OpenStudy (anonymous):

Write the equation of the line that passes through each pair of points in slope-intercept form. (-1,7),(4,-3) P L E A S E H E L P .

OpenStudy (anonymous):

Tell me this first: How do you find your slope?

OpenStudy (anonymous):

Slope: (y2-y1)/(x2-x1)

sam (.sam.):

(y2-y1)/(x2-x1) find your m,or slope

OpenStudy (anonymous):

This Is How You Find Your Slope. y=mx+b

OpenStudy (anonymous):

No Umme, your slope is the m-value. m=(y2-y1)/(x2-x1)

OpenStudy (anonymous):

No that's the equation of the line, not slope.

sam (.sam.):

\[\huge Slope=\frac{y _{2}-y _{1}}{x _{2}-x _{1}}\]

OpenStudy (anonymous):

Umme, tell me what you got for your slope and I'll tell you more about the problem.

OpenStudy (anonymous):

IGot y=-2x-11 For My Answer.

OpenStudy (anonymous):

m = (-3-7)/(4-(-1)) m= -10/5 m=-2/1 m=-2

OpenStudy (anonymous):

Okay This Is What IDid.

OpenStudy (anonymous):

y2-y1=15-5 x2-x1=-8-2 = 10/10=1 y=mx+b 5=-1(2)+b 5=-2+b

OpenStudy (anonymous):

5=-2+b +-2 +-2 3=b y=-1x-3

OpenStudy (anonymous):

Read through what I typed very carefully.

OpenStudy (anonymous):

For a second I thought you hit and run lol

OpenStudy (anonymous):

How Did You Get 4 Though.?

OpenStudy (anonymous):

&Lol.:P No.

OpenStudy (anonymous):

Oh man am I not here tonight!

OpenStudy (anonymous):

SORRY! Tiny little error, I will repost.

OpenStudy (anonymous):

It's Okay. :P &Thanks.

OpenStudy (anonymous):

Okay, now you have 2 things: Your point: You can use either (-1,7) or (4,-3), doesn't matter which one you choose. Your slope: m=-2 Now you need to find your b-value. y=mx+b I'll use (-1,7) for (x,y) and m=-2 7=-2(-1) +b b = 5 Now this is the easy part. All you need for slope-intercept form is your slope and your b-value. y=mx+b y=-2x +5

OpenStudy (anonymous):

&How'd You Get 5.?:\ I'm Confused.

OpenStudy (anonymous):

Doesn't Negative 2 And Negative One Equal To Positive 3.?

OpenStudy (anonymous):

7 = -2(-1) +b 7 = 2 + b 7-2 = b 5 = b

OpenStudy (anonymous):

Oh.Okay.Nevermind.My Mistake. :P

OpenStudy (anonymous):

2*1 = 2 2*(-1) = -2 (-2)(-1) = 2 :-)

OpenStudy (anonymous):

No problemo

OpenStudy (anonymous):

Thanks A Ton. :D

OpenStudy (anonymous):

&:D

OpenStudy (anonymous):

So, in summary: You are given 2 points. To find your line, follow these steps: 1: Find the slope 2. Find your b-value with y=mx+b (plug in a point, either point will work) and your m-value. Solve for b. 3. You'll have your m-value and your b-value. Put ONLY THOSE values into y=mx+b And there you have it, your line!

OpenStudy (anonymous):

Thanks Again. :)

OpenStudy (anonymous):

No problem :-) I would suggest Paul's online notes for any of your algebra needs. And calculus when you take it in college.

OpenStudy (anonymous):

What's Paul's Online Notes.? ._.

OpenStudy (anonymous):

http://tutorial.math.lamar.edu/Classes/Alg/Lines.aspx Paul = math God ...imho :-)

OpenStudy (anonymous):

Oh. Okay. Thanks.

OpenStudy (anonymous):

&Another Thing, For This Same Thing, IHave Another Problem. It's (2,5),(-8,15)

OpenStudy (anonymous):

The Answer IGot Is y= -1x-3.

OpenStudy (anonymous):

Is It Right.?

OpenStudy (anonymous):

lemme try... one sec

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

I should really do these on paper :-(

OpenStudy (anonymous):

True. But Oh Well. :P

OpenStudy (anonymous):

Is My Answer Right Though.?

OpenStudy (anonymous):

So I got m= -1 and b=7

OpenStudy (anonymous):

5 = (-1)(2) +b 5 = -2 +b 7 =b

OpenStudy (anonymous):

If you put in both of your points, this will give you a step-by-step walk through on how to do these: http://www.webmath.com/equline1.html

OpenStudy (anonymous):

Okay. Thanks.

OpenStudy (anonymous):

Whoahh.! This Website's Really Good.!

OpenStudy (anonymous):

There is another formula for doing it all in one step. Question is, if you can remember the formula. http://www.youtube.com/watch?v=GFM8NOe_XM4

OpenStudy (anonymous):

Thank You.

OpenStudy (anonymous):

IDon't Know How They Got 7 Though.

OpenStudy (anonymous):

5 = (-1)(2) +b 5 = -2 +b 7 =b

OpenStudy (anonymous):

y=mx+b y=5 m=-1 x=2

OpenStudy (anonymous):

Oh.Okay.Thanks Again.

OpenStudy (anonymous):

Welcome :-). Keep working at it. Do as many of these problems as possible. That's how these things work... you bang your head against the wall until something finally sticks :-D

OpenStudy (anonymous):

Good luck!

OpenStudy (anonymous):

So 5+2 = 7 Or What.? o.O

OpenStudy (anonymous):

Yep